OpenStudy (anonymous):

solve f(x)+g(x)=6 f(x) = log2(x-7) and g(x) = log2(3x-5)

OpenStudy (anonymous):

log2(x-7)(3x-5)=6 (x-7)(3x-5)=36... you can take it from here

OpenStudy (anonymous):

sure. why not. \[f(x) + g(x) = 6\] \[\log _{2}(x-7) + \log_{2}(3x-5) = 6\] \[\log_{2}(x-7)(3x-5)=6\]

OpenStudy (anonymous):

...you don't square both sides. In logarithms, its log based 2. Therefore you raise the left and right side of the equation as a power of 2. so, it is 2 ^ 6 = (x-7)(3x-5)

OpenStudy (anonymous):

oh and i forgot to mention that my only answers are a) 9 b) -9 c) 512 d) -512

OpenStudy (anonymous):

then plug them in and play. find out which one works....